# Semilinear spaces of matrices with transitivity, rank or spectral constraints

### v1

@@ 1.5 The strategy, and intransitive operator spaces
Let $U$ and $V$ be right-$\mathbb{D}$-vector spaces and $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$. We say that $\mathcal{S}$ is weakly primitively intransitive whenever there is no $1$-dimensional $\mathbb{D}$-linear subspace $V^{\prime}$ of $V$ such that, for the standard projection $\pi:V\twoheadrightarrow V/V^{\prime}$, the operator space $\pi\mathcal{S}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V/V^{\prime})$ satisfies $\operatorname{\operatorname{trk}}(\pi\mathcal{S})\leq\operatorname{\operatorname{tr}}(\mathcal{S})-d$.
@@
Let $U$ and $V$ be right-$\mathbb{D}$-vector spaces and $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$. We say that $\mathcal{S}$ is primitively intransitive whenever it is intransitive and there is no nonzero linear subspace $V^{\prime}$ of $V$ such that, for the standard projection $\pi:V\twoheadrightarrow V/V^{\prime}$, the operator space $\pi\mathcal{S}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V/V^{\prime})$ is intransitive.
@@
Let $U$ and $V$ be right-$\mathbb{D}$-vector spaces, $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$, and assume that there exists a $1$-dimensional $\mathbb{D}$-linear subspace $V^{\prime}$ of $V$ such that, for the standard projection $\pi:V\twoheadrightarrow V/V^{\prime}$, the operator space $\pi\mathcal{S}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V/V^{\prime})$ satisfies $\operatorname{\operatorname{trk}}(\pi\mathcal{S})\leq\operatorname{\operatorname{tr}}(\mathcal{S})-d$. If $\mathcal{S}$ is intransitive, then we deduce that $\operatorname{\operatorname{trk}}(\pi\mathcal{S})<\dim_{\mathbb{F}}(V/V^{\prime})$ and hence $\pi\mathcal{S}$ is intransitive as well: this shows that $\mathcal{S}$ is not primitively intransitive. ∎
A key insight of Atkinson and Lloyd can be rephrased in saying that, provided that the field $\mathbb{F}$ has enough elements, every weakly primitively intransitive space is deeply intransitive. The primitively intransitivity condition has the shortcoming that it is generally not inherited by $\mathcal{S}^{V^{\prime}}$ when $V^{\prime}$ is a linear subspace of $V$, in contrast with the deep intransitivity condition. However, it is the most convenient condition when applying the previous theorem. Indeed, we shall prove – under the usual cardinality conditions on $\mathbb{F}$ – that every irreducible optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ is primitively intransitive, whereas we are not aware of any direct proof that it must be deeply intransitive.
@@ 1.6 Structure of the article
As hinted at in the previous paragraph, our focus is now shifting from the trivial spectrum subspaces, which are the prime motivation for this article, to the deeply intransitive operator spaces, which can be viewed as a more general topic. We expect that studying the later will yield many applications in the future, just like Atkinson’s theorem has found many unexpected applications in the recent past (see e.g. [19, 25]). In an early version of the present article, we wrote a proof of the result on trivial spectrum subspaces alone. Surprising as it may seem, we found later that the result on deeply intransitive operator spaces is easier to prove. The icing on the cake is that it also holds more potential for future applications.
@@
The next section (Section 4) introduces the key technical tools to study the intransitive operator spaces: there we make no claims of originality, as we will use the generic matrix machinery that allowed Atkinson to prove his theorem, a machinery which has already demonstrated its remarkable efficiency for such problems. The main results on intransitive operator spaces are obtained in Section 5: there, the technical details for a general division ring $\mathbb{D}$ are substantially more difficult than in Atkinson’s article.
@@
The remaining sections are devoted to applications of trivial spectrum subspaces. In Section 7, we obtain the greatest possible dimension for an $\mathbb{F}$-affine subspace of $\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})$ with rank at least $r$ (Section 7.1), and in Section 7.2 we characterize the spaces with the greatest possible dimension when $n=p=r$ (i.e., affine subspaces of nonsingular matrices). The characterization for a general $r$ is very difficult (see [16] for the case $d=1$, which takes no less than twenty-three pages of proof), so we believe the reader will be relieved to learn that this issue is postponed to a subsequent article.
The last section is devoted to the applications of $\mathbb{F}$-trivial spectrum subspaces to subspaces of $\mathbb{F}$-diagonalisable matrices. There, we successfully adapt the method featured in [23], and we also give a brand new result for real vector spaces of complex diagonalisable matrices. This new application involves the main theorem on deeply intransitive operator spaces in a case where the space under consideration has its dimension slightly lower than the critical one. We finish with an adaptation to spaces of semi-simple endomorphisms and with considerations on spaces of endomorphisms in which only the zero endomorphism is nilpotent.
@@ 2 On $\mathbb{D}$-vector spaces seen as $\mathbb{F}$-vector spaces
2.1 Traces
Throughout, we will need to consider $\mathbb{F}$-linear forms on $\mathbb{D}$. There is a canonical choice of such a linear form: the trace of $\mathbb{D}$ over $\mathbb{F}$, denote by $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}$ (with capital T), maps every $a\in\mathbb{D}$ to the trace of the $\mathbb{F}$-endomorphism $x\mapsto ax$ of $\mathbb{D}$. Consequently, it is an $\mathbb{F}$-linear form that enjoys the property that $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}(ab)=\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}(ba)$ for all $a,b$ in $\mathbb{D}$. Let $a\in\mathbb{D}$. Denoting by $\pi=t^{n}-\alpha t^{n-1}+\cdots$ the minimal polynomial of $a$ in the $\mathbb{F}$-algebra $\mathbb{D}$, it is easily seen that $n$ divides $d$ and that $x\mapsto ax$ is the direct sum of $\frac{d}{n}$ cyclic endomorphisms of $\mathbb{F}$-vector spaces with minimal polynomial $\pi$, which yields $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}(a)=\frac{d}{n}\,\alpha.$ Applying the same line of reasoning to $x\mapsto xa$ yields that $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}(a)$ is also the trace of the $\mathbb{F}$-linear endomorphism $x\mapsto xa$. Moreover $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}(\varphi(a))=\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}(a)$ for every automorphism $\varphi$ of the $\mathbb{F}$-algebra $\mathbb{D}$.
Alas, in some cases $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}$ might be useless. This is the case for example if $\mathbb{D}$ is a totally radicial extension of $\mathbb{F}$: in that case $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}$ is identically zero. The same holds if $\mathbb{D}$ is a quaternion algebra and $\operatorname{\operatorname{char}}(\mathbb{F})=2$: in that case the minimal polynomial of every element of $\mathbb{D}$ has degree $1$ or $2$, and hence the trace of such an element is either $2$ times or $4$ times the trace of this polynomial, so it is zero in any case. If however $\operatorname{\operatorname{char}}(\mathbb{F})=0$ then $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}(1_{\mathbb{D}})=d\,1_{\mathbb{F}}$ is nonzero and we can safely use $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}$ as a nonzero linear form on $\mathbb{D}$.
@@ 2.4 Multiplicative quadratic forms on division algebras
Our study will require the following result. It can be extracted from the general theory of composition algebras, but since we are in the very specific context of a division ring $\mathbb{D}$, the proof can be considerably shortened. We include a proof below for non-specialists, but advise the reader to skip it at first reading.
@@
Proof.
Throughout we consider the polar form $b:(x,y)\in\mathbb{D}^{2}\mapsto q(x+y)-q(x)-q(y)$ of $q$, and note that the multiplicativity of $q$ yields
$\forall(a,x,y)\in\mathbb{D}^{3},\;b(ax,ay)=q(a)\,b(x,y)=b(xa,ya).$ (2)
Assume first that $q$ is degenerate. Then $\operatorname{\operatorname{char}}(\mathbb{F})=2$. We start by proving that $q$ is totally degenerate (i.e., additive). Let us consider the radical $R$ of $b$. We have $R\neq\{0\}$ by assumption. Let us choose $x\in R\smallsetminus\{0\}$. For all $a\in\mathbb{D}\smallsetminus\{0\}$ we have
$\forall y\in\mathbb{D},\;b(ax,ay)=q(a)\,b(x,y)=0,$
and since $\{ay\mid y\in\mathbb{D}\}=\mathbb{D}$ we deduce that $ax\in R$. Hence $R=\mathbb{D}$, meaning that $q$ is additive.
Next, for all $x,y$ in $\mathbb{D}$ we find
q(xyyx)=q(xy)q(yx)=q(x)q(y)q(y)q(x)=0,q(xy-yx)=q(xy)-q(yx)=q(x)q(y)-q(y)q(x)=0,
and hence xyyx=0xy-yx=0 because $q$ is nonisotropic. Hence $\mathbb{D}$ is commutative.
Let $x\in\mathbb{D}$. Then
$q(x^{2}-q(x))=q(x^{2})-q(q(x).1_{\mathbb{D}})=q(x)^{2}-q(x)^{2}q(1_{\mathbb{D}})=0,$
and hence $x^{2}=q(x)\in\mathbb{F}$ because $q$ is nonisotropic. Therefore $\mathbb{D}$ is a hyper-radicial extension of $\mathbb{F}$, and $q(x)=x\sigma(x)$ for all $x\in\mathbb{D}$, where $\sigma=\operatorname{\operatorname{id}}_{\mathbb{D}}$ is the sole involution of the $\mathbb{F}$-algebra $\mathbb{D}$.
Now, we deal with the case where $q$ is nondegenerate. We introduce the trace map $\operatorname{\operatorname{tr}}:x\in\mathbb{D}\mapsto b(1,x)\in\mathbb{F}$ and note that $\operatorname{\operatorname{tr}}(1)=2$. The key point consists in proving the quadratic identity
$\forall x\in\mathbb{D},\;x^{2}-\operatorname{\operatorname{tr}}(x)x+q(x)=0.$ (3)
To do so, we take arbitrary elements $y$ and $z$ in $\mathbb{D}$ and explore the identity $q(x+y)q(x+z)=q((x+y)(x+z))=q(x^{2}+yx+xz+yz)$. Thanks to $b$, we expand
$q(x^{2}+yx+xz+yz)=\underbrace{q(x^{2})}_{a_{4}}+\underbrace{b(x^{2},yx+xz)}_{a_{3}}+\underbrace{b(x^{2},yz)+q(yx+xz)}_{a_{2}}\\ +\underbrace{b(yx+xz,yz)}_{a_{1}}+\underbrace{q(yz)}_{a_{0}}$
and
$q(x+y)q(x+z)=\underbrace{q(x)^{2}}_{b_{4}}+\underbrace{q(x)(b(x,y)+b(x,z))}_{b_{3}}+\underbrace{q(x)(q(y)+q(z))+b(x,y)b(x,z)}_{b_{2}}\\ +\underbrace{q(z)b(x,y)+q(y)b(x,z)}_{b_{1}}+\underbrace{q(y)q(z)}_{b_{0}}.$
It is easily checked, by using either the multiplicativity of $q$ or (2), that $b_{i}=a_{i}$ for all $i\in\{0,1,3,4\}$, which leaves us with
$q(x)(q(y)+q(z))+b(x,y)\,b(x,z)=b(x^{2},yz)+q(yx+xz),$
and since q(yx+xz)=q(yx)+q(xz)+b(yx,xz)=q(y)q(x)+q(x)q(z)+b(yx,xz)q(yx+xz)=q(yx)+q(xz)+b(yx,xz)=q(y)q(x)+q(x)q(z)+b(yx,xz), this is further simplified as
$b(x,y)\,b(x,z)=b(x^{2},yz)+b(yx,xz).$
From there, we apply this identity to $y=1$ and an arbitrary $z$ to find $b(q(x),z)=q(x)\,b(1,z)=b(x,xz)=b(yx,xz)=b(x,1)\,b(x,z)-b(x^{2},z)=b(\operatorname{\operatorname{tr}}(x)x-x^{2},z)$, which yields $q(x)=\operatorname{\operatorname{tr}}(x)x-x^{2}$ because $q$ is nondegenerate. Next, we introduce the adjunction mapping
$x\in\mathbb{D}\mapsto x^{\star}:=\operatorname{\operatorname{tr}}(x)-x\in\mathbb{D},$
which is $\mathbb{F}$-linear and clearly satisfies $\operatorname{\operatorname{tr}}(x^{\star})=\operatorname{\operatorname{tr}}(x)$ and $(x^{\star})^{\star}=x$ for all $x\in\mathbb{D}$. We observe that (3) can be reformulated into
$\forall x\in\mathbb{D},\;q(x)=xx^{\star}=x^{\star}x.$
In particular, for all $x,y$ in $\mathbb{D}^{\times}$, we have $(xy)(xy)^{\star}=q(xy)=q(x)q(y)=xq(y)x^{\star}=xyy^{\star}x^{\star}$ and hence $(xy)^{\star}=y^{\star}x^{\star}$. It follows that $(-)^{\star}$ is an involution of the $\mathbb{F}$-algebra $\mathbb{D}$.
Assume now that $d>1$. Since $\mathbb{D}$ is a division ring, we deduce from (3) that it has a $2$-dimensional subfield, and hence $d$ is even. Assume that $d=2$. Since $q$ is nondegenerate, the trace does not vanish, so by linearity it does not vanish on $\mathbb{D}\smallsetminus\mathbb{F}$, and we deduce from (3) that $\mathbb{D}$ is a separable extension of $\mathbb{F}$ (by choosing $x\in\mathbb{D}\smallsetminus\mathbb{F}$ such that $\operatorname{\operatorname{tr}}(x)\neq 0$). And clearly $(-)^{\star}$ is not the identity, so it is the standard involution of $\mathbb{D}$ over $\mathbb{F}$.
Assume finally that $d>2$, and set p:=d2p:=d-2. Since $q$ is nondegenerate, the radical of its restriction to $H:=\operatorname{\operatorname{Ker}}(\operatorname{\operatorname{tr}})$ has dimension at most $1$, so we can find in $H$ an $\mathbb{F}$-linear subspace $V$ of dimension $p$ on which $q$ remains nondegenerate. Then the identity $\forall x\in V,\;x^{2}=-q(x)$ yields a (unique) homomorphism of $\mathbb{F}$-algebras $\Phi:\mathcal{C}(V,-q_{|V})\rightarrow\mathbb{D}$ such that $\forall x\in V,\;\Phi(x)=x$, where $\mathcal{C}(V,-q_{|V})$ stands for the Clifford algebra of the regular quadratic space $(V,-q_{|V})$. Classically, $\mathcal{C}(V,-q_{|V})$ is a simple algebra with dimension $2^{p}$ over $\mathbb{F}$, because $p$ is even. Hence $\Phi$ is injective, which leads to $2^{p}\leq d$ and hence to $d=4$. Then $\Phi$ is an isomorphism and $\mathbb{F}\cap V=\{0\}$, and we have recognized from $\Phi$ that $\mathbb{D}$ is a quaternion algebra over $\mathbb{F}$ and that the quaternionic conjugation $x\mapsto\overline{x}$ inherited from $\Phi$ satisfies $x\overline{x}=q(x)$ for all $x\in V$, which yields that it coincides with $(-)^{\star}$ on $V$. Hence $x\mapsto x^{\star}$ and $x\mapsto\overline{x}$ coincide on the $\mathbb{F}$-subalgebra generated by $\mathbb{F}\oplus V$, which must be $\mathbb{D}$ because it is a division algebra over $\mathbb{F}$ with dimension $3$ or $4$ that contains a subfield with dimension $2$ over $\mathbb{F}$. Therefore $x\mapsto x^{\star}$ is the quaternionic conjugation of $\mathbb{D}$, and our proof is completed. ∎

### v2

@@ 1.5 The strategy, and intransitive operator spaces
Let $U$ and $V$ be $\mathbb{D}$-vector spaces and $S$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$. We say that $S$ is weakly primitively intransitive whenever there is no $1$-dimensional $\mathbb{D}$-linear subspace $V^{\prime}$ of $V$ such that, for the standard projection $\pi:V\twoheadrightarrow V/V^{\prime}$, the operator space $\pi S\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V/V^{\prime})$ satisfies $\operatorname{\operatorname{trk}}(\pi S)\leq\operatorname{\operatorname{tr}}(S)-d$.
@@
Let $U$ and $V$ be $\mathbb{D}$-vector spaces and $S$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$. We say that $S$ is primitively intransitive whenever it is intransitive and there is no nonzero linear subspace $V^{\prime}$ of $V$ such that, for the standard projection $\pi:V\twoheadrightarrow V/V^{\prime}$, the operator space $\pi S\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V/V^{\prime})$ is intransitive.
@@
Let $U$ and $V$ be $\mathbb{D}$-vector spaces, $S$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$, and assume that there exists a $1$-dimensional $\mathbb{D}$-linear subspace $V^{\prime}$ of $V$ such that, for the standard projection $\pi:V\twoheadrightarrow V/V^{\prime}$, the operator space $\pi S\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V/V^{\prime})$ satisfies $\operatorname{\operatorname{trk}}(\pi S)\leq\operatorname{\operatorname{tr}}(S)-d$. If $S$ is intransitive, then we deduce that $\operatorname{\operatorname{trk}}(\pi S)<\dim_{\mathbb{F}}(V/V^{\prime})$ and hence $\pi S$ is intransitive as well: this shows that $S$ is not primitively intransitive. ∎
A key insight of Atkinson and Lloyd can be rephrased in saying that, provided that the field $\mathbb{F}$ has enough elements, every weakly primitively intransitive space is deeply intransitive (in their terminology, this is the statement that every matrix space that satisfies condition (C) has the row property). The primitive intransitivity condition has the shortcoming that it is generally not inherited by $S^{V^{\prime}}$ when $V^{\prime}$ is a linear subspace of $V$, in contrast with the deep intransitivity condition. However, it is the most convenient condition when applying the previous theorem. Indeed, we shall prove – under the usual cardinality conditions on $\mathbb{F}$ – that every irreducible optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ is primitively intransitive, whereas we are not aware of any direct proof that it must be deeply intransitive.
@@ 1.6 Structure of the article
As hinted at in the previous paragraph, our focus, which had already shifted from unital affine subspaces to trivial spectrum linear subspaces, is now shifting from trivial spectrum subspaces to deeply intransitive operator spaces, a more profound and general problem. We expect that studying the later will yield many applications in the future, just like Atkinson’s theorem has found many unexpected applications in the recent past (see e.g., [13, 19]). In an early version of the present article, we wrote a proof of the result on trivial spectrum subspaces alone. Surprising as it may seem, we found later that the result on deeply intransitive operator spaces was much easier to prove. The icing on the cake is that it also holds far more potential for future applications.
@@
The next section (Section 4) introduces the key technical tools to study the intransitive operator spaces: there we make no claims of originality, as we will use the generic matrix machinery that allowed Atkinson to prove his theorem, a machinery which has already demonstrated its remarkable efficiency for such problems. The main results on intransitive operator spaces are obtained in Section 5: there, the technical details for a general division ring $\mathbb{D}$ are much more difficult than in Atkinson’s article.
@@
Finally, in Section 7 we conclude the study by solving the initial problem of classifying the optimal unital $\mathbb{F}$-affine subspaces of a simple $\mathbb{F}$-algebra, up to equivalence.
1.7 Potential applications
There are many applications of the results featured in this article, but due to a lack of space we have decided to keep them out of this article. However, we can quickly discuss two of them.
The first application is towards affine subspaces of operators with rank bounded below. In this, one takes finite-dimensional $\mathbb{D}$-vector spaces $U$ and $V$, together with a positive integer $r\leq\min(\dim_{\mathbb{D}}U,\dim_{\mathbb{D}}V)$, and one asks about the greatest possible dimension for an $\mathbb{F}$-affine subspace of elements of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ in which every operator has rank at least $r$, and how to classify the spaces of greatest possible dimension. The special case $r=\dim_{\mathbb{D}}U=\dim_{\mathbb{D}}V$ is essentially the problem of unital affine subspaces in the simple $\mathbb{F}$-algebra $\operatorname{\operatorname{End}}_{\mathbb{D}}(U)$. Deriving the greatest possible dimension in the general case is not difficult, but classifying the optimal spaces is highly non-trivial: see [11] for the special case where $\mathbb{D}=\mathbb{F}$.
A second potential application is towards $\mathbb{F}$-linear subspaces of $\mathbb{F}$-diagonalisable elements of a simple $\mathbb{F}$-algebra $\mathcal{A}$. There, an element $a\in\mathcal{A}$ is called $\mathbb{F}$-diagonalisable whenever its minimal polynomial over $\mathbb{F}$ splits with simple roots, or equivalently when $x\mapsto ax$ is a diagonalisable endomorphism of the $\mathbb{F}$-vector space $\mathcal{A}$. When $\mathcal{A}=\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ this problem has already been studied, in particular over the field of real numbers: for an arbitrary field $\mathbb{F}$, any $\mathbb{F}$-linear subspace of $\mathbb{F}$-diagonalisable elements of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ has dimension at most $\frac{n(n+1)}{2}$, with equality attained only for conjugates of the space $\operatorname{\operatorname{S}}_{n}(\mathbb{F})$ of all symmetric matrices (and if $n\geq 2$ such spaces exist only if $\mathbb{F}$ is formally real and Pythagorean). A connection with the problem of trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ was discovered and exploited in [18]: In a sequel to this article we will use this connection to generalize the study of linear subspaces of $\mathbb{F}$-diagonalisable elements in any simple $\mathbb{F}$-algebra, and in particular we will obtain the structure of the $\mathbb{R}$-linear subspaces of $\mathbb{R}$-diagonalisable matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{C})$: it will turn out that the optimal spaces are the conjugates of the space $\operatorname{\operatorname{H}}_{n}(\mathbb{C})$ of all Hermitian matrices. There is also a quaternionic analogue of this result.
@@ 2 On $\mathbb{D}$-vector spaces seen as $\mathbb{F}$-vector spaces
From now on, $\mathbb{D}$ denotes a division algebra that contains $\mathbb{F}$ as a central subfield, and we denote by $d$ the dimension of $\mathbb{D}$ over $\mathbb{F}$.
@@ 2.3 Multiplicative quadratic forms on division algebras
Our study will require the following classical result from the theory of composition algebras.
@@
The result is rarely presented in this way in the literature, in particular because most references on the topic do not restrict the scope to nonisotropic quadratic forms. Hence, some comments are necessary. If $q$ is non-degenerate then the theory of unital composition algebras yields that either $\mathbb{D}$ is a Cayley algebra over $\mathbb{F}$ or $\mathbb{D}$ is of separable quadratic type over $\mathbb{F}$ and $q:x\mapsto x\sigma(x)$ for the standard involution $\sigma$ of $\mathbb{D}$ over $\mathbb{F}$ (see theorem 33.17 in [7]). Yet, here the multiplication map is associative, so Cayley algebras are discarded. Finally, if $q$ is degenerate then $\operatorname{\operatorname{char}}(\mathbb{F})=2$ and it is known in that case that the radical of $q$ equals the full space $\mathbb{D}$ (see proposition 33.4 in [7]), and from there it is known that $\mathbb{D}$ is a hyper-radicial extension of $\mathbb{F}$ and that $q:x\mapsto x^{2}$ (see corollary 33.6 in [7]).